OpenStudy (anonymous):

PLEASE SHOW WORK, SO I UNDERSTAND!! Gino designed a garden for his rectangular backyard. The design he drew on paper is 7 inches wide. The width of his backyard is 42 feet. If his drawing is to scale and is 8 inches long, how many feet long is his backyard? Write a proportion using words. Then, use the proportion to calculate the yard length. Show your work.

OpenStudy (anonymous):

7/42 =8/X 7X=336 336/7 = 48 ft long